OpenStudy (anonymous):

For which system of equations is (–2, 2) the solution? A. 3x – 5y = –16 –2x + 3y = 10 B. –4x + 2y = 6 2x – 3y = 10 C. 4x + 3y = 2 5x + 4y = –2 D. 6x + 5y = –2 –8x + 5y = –6

OpenStudy (whpalmer4):

It's the solution for the system of equations where both equations are true if \(x=-2\) and \( y=2\).

OpenStudy (whpalmer4):

For example, I'll test answer choice D: \[6x+5y=-2\]\[-8x+5y=-6\] \[6(-2)+5(2) = -2\]\[-12+10=-2\]So far so good, but we have to test both equations to be sure. \[-8(-2)+5(2) = -6\]\[16+10=-6\]Oops. D is not the correct answer, because one of the equations is not true when \((-2,2)\) is substituted for \((x,y)\).